OpenStudy (anonymous):

how do I find the derivative of f(x)= sin[ln(cosx^3)]

OpenStudy (ash2326):

Use chain rule here \[\frac{d}{dx} f( g(x))=f'(x)\times \frac{d}{dx} g(x)\] Could you try ?

OpenStudy (anonymous):

would it be cos[lncosx^3] X (-sinx^3)/(cosx^3) ?

OpenStudy (ash2326):

times the derivative of x^3 \[\cos (\ln{\cos x^3})\times \frac{1}{\cos x^3}\times (-\sin x^3)\times \frac{d}{dx} (x^3)\] Good work btw

OpenStudy (anonymous):

Thanks !

OpenStudy (ash2326):

You're welcome :D